OpenStudy (anonymous):

H1P2: DECOMPOSITION OF AMMONIUM NITRATE

OpenStudy (anonymous):

Solid NH4NO3 (ammonium nitrate) decomposes on heating to 400°C, forming N2O gas and water vapor, H2O. (a)Write a balanced chemical equation. incorrect (b) Calculate the number of grams of H2O that will form on decomposition of 0.10 mole of ammonium nitrate.

OpenStudy (anonymous):

so first write the balanced equation, ill write it you balance it: NH4NO3 -> N2O + H2O and then from the reaction ratio you calculate mass of H2O produced from 0,1 mole of NH4NO3 n(H2O)/n(NH4NO3) = ?/? where ?/? is ratio from balanced chemical reaction and then n(H2O) = ?/? * n(NH4NO3) and n=m/M so m(H2O)= ?/? * n(NH4NO3) * M(H2O) any questions?

OpenStudy (anonymous):

i have balanced the equation but im still uncertain how we make work out how many grams of h20 we can produce from 0.10 moles of amonium nitrate

OpenStudy (anonymous):

NH4NO3 -> N2O + 2H2O -------------------------- n(H2O)/n(NH4NO3) = 2/1 = 2 so n(H2O) = 2 * n(NH4NO3) n(H2O) = 2*0,1 mol n(H2O) = 0,2 mol n=m/M m(H2O) = n(H2O) * M(H2O) m(H2O) = 0,2 mol * 18 g mol-1 m(H2O) = 3,6 g